OpenStudy (anonymous):

Taylor Stiles owns 75 shares of a Strong Pharmaceutical Company stock (SPC). After the quarterly dividends are reinvested, Taylor owns 78 shares. SPC is selling at $45.50 a share. What is the annual yield for SPC?

OpenStudy (anonymous):

annual yield= annual dividend per share / cost per share

OpenStudy (anonymous):

omg marry me! :DDD

OpenStudy (anonymous):

Initially you have 75 shares. After the first Quarter, after 3 months, you now have 3 more shares than you initially had because of the reinvested dividends so if we multiply that how much each share is: \[3 shares \times \frac{ 45.50$ }{ share }= 136.5$\] So now we know how much we made off those 75 shares but how much did we make individually per share? To do that we set up another ratio to find the amount made per share now that we know how much we make per 75 shares quarterly: \[\frac{ 136.5$ }{ 75 shares } = 1.82$/ share\]

OpenStudy (anonymous):

Oh thank you!!!

OpenStudy (anonymous):

That's exactly what I needed help with

OpenStudy (anonymous):

Now that you have figured out how much you make individually per share you should be able to find how much you make in the second, third, and fourth quarter and how much shares can be bought and how much in dividends you will have at the end of the 4 quarters.

OpenStudy (anonymous):

My pleasure.

OpenStudy (anonymous):

so 16% ? because you multiply .04 x 4 (because it's quarterly,correct) ? @Johnbc

OpenStudy (anonymous):

The dividends are reinvested so you first had 75 and then 78 so you need to find out how much you make in the seconds and how many shares you can buy with that amount made and add them to the 78 and then using your new total shares do the same for the third and repeat for the 4th and find out the total dividends in those 4 quarters

OpenStudy (anonymous):

got it!! thank you:)

OpenStudy (anonymous):

Awesome!, my pleasure.